THE FOURIER TRANSFORM MICROWAVE (FTMW) SPECTRA OF CYCLOHEXENE OXIDE AND ITS ARGON COMPLEX

\underline{DANIEL J. FROHMAN}, STEWART E. NOVICK, and WALLACE C. PRINGLE, Department of Chemistry, Wesleyan University, 52 Lawn Ave., Middletown, CT 06459-0180.

The microwave spectrum of cyclohexene oxide and its isotopologues have been observed and assigned, improving upon previous rotational studies of this molecule.^{a}^{,}^{b} Additionally, the ^{17}O isotopomer of cyclohexene oxide and the Ar complex of the normal isotopologue of cyclohexene oxide have been fit for the first time. Fits for the ^{13}C-cyclohexene oxide Ar complexes will also be presented.

aTatsuya Ikeda, Roger Kewley, and R. F. Curl, Jr. J. Mol. Spectrosc., \underline{44} (1972), 459-469.

bRaquel Sánchez, Susana Blanco, Juan C. López, and José L. Alonso. J. Mol. Struct., \underline{780-781} (2006), 57-64.